# XLand-100B: A Large-Scale Multi-Task Dataset for In-Context Reinforcement Learning

### v1

@@ Abstract
Following the success of the in-context learning paradigm in large-scale language and computer vision models, the recently emerging field of in-context reinforcement learning is experiencing a rapid growth. However, its development has been held back by the lack of challenging benchmarks, as all the experiments have been carried out in simple environments and on small-scale datasets. We present XLand-100B, a large-scale dataset for in-context reinforcement learning based on the XLand-MiniGrid environment, as a first step to alleviate this problem. It contains complete learning histories for nearly $30,000$ different tasks, covering $100$B transitions and $2.5$B episodes. It took $50,000$ GPU hours to collect the dataset, which is beyond the reach of most academic labs. Along with the dataset, we provide the utilities to reproduce or expand it even further. With this substantial effort, we aim to democratize research in the rapidly growing field of in-context reinforcement learning and provide a solid foundation for further scaling. The code is open-source and available under Apache 2.0 licence at https://github.com/dunno-lab/xland-minigrid-datasets.
@@ 1 Introduction
However, despite the rapid adoption of the transformer architecture in reinforcement learning (RL) after the release of Decision Transformer (DT) (Chen et al., 2021; Hu et al., 2022; Agarwal et al., 2023; Li et al., 2023), models with in-context learning capabilities appeared only recently. This delay is caused by a number of reasons. Firstly, to transition from in-weights to in-context learning, a model should be trained on tens of thousands of unique tasks (Kirsch et al., 2022). Unfortunately, even the largest RL datasets currently contain only hundreds or tasks (Padalkar et al., 2023). Secondly, it was necessary to determine the right way to provide a context to a transformer and develop the data collection pipeline, which for many methods (Laskin et al., 2022; Lee et al., 2023; Shi et al., 2024) was different from what is commonly available in existing datasets (see Section 3).
Due to the lack of suitable datasets and the high cost of collecting data in existing environments, the recent wave of in-context RL research (Laskin et al., 2022; Lee et al., 2023; Norman & Clune, 2023; Kirsch et al., 2023; Sinii et al., 2023; Zisman et al., 2023) used environments with very simple task distributions, where it was feasible to collect datasets with hundreds of tasks. While these benchmarks are affordable, they are not suitable for the comparison of methods at scale on tasks with high diversity and difficulty, which is essential for real-world applications. Because of this, the development of in-context RL is currently hindered by these factors. We believe it is crucial to address these barriers, given the essential role of in-context learning in the path to foundation models and truly generalist agents (Team et al., 2021, 2023; Kirsch et al., 2023; Lu et al., 2024; Liu et al., 2024).
@@
Data # tasks # transitions # episodes Size Open-Source Enables ICL XLand-100B (ours) 28,876 100B 2.5B 320GB JAT 157 323M N/A 1TB GATO 596 1.5T 63M N/A Open X-Embodiment 527 N/A 2.4M 9TB AlphaStar Unplugged 1 21B 2.8M N/A ? NetHack 1 3.5B 110K 97 GB D4RL 11 N/A 40 M 4.8GB V-D4RL 4 2.4M N/A 17GB D5RL 50 N/A 2500 N/A RL Unplugged 90 80M N/A 54.1TB
@@ 2.2 XLand-MiniGrid
Figure 1: Visualization of a generic XLand-MiniGrid environment. Grid layout should be selected in advance, while the positions of the objects are randomized on each reset.
@@
Environment. XLand-MiniGrid is a complete rewrite of MiniGrid (Chevalier-Boisvert et al., 2023) in JAX (Bradbury et al., 2018), incorporating a notion of rules and goals from XLand (Team et al., 2023). Leveraging JAX, it can run on a GPU or TPU accelerators at millions steps per seconds. At its core, it is a goal-oriented grid-world environment with simple underlying dynamics, partial observability and sparse rewards. The action space is simple, consisting mainly of navigation and interaction with game objects, such as opening a door or picking and placing items. Observations are symbolic "images" encoding the agent surrounding as tile and color ID’s. Rules are functions that can change the state of the environment based on some conditions, e.g. when two specific objects are places near each other, they both disappear and one new object is placed. Goals are similar, except they only validate a predefined conditions and do not change anything. Composing different rules and goals together we can create new tasks with varying reward and dynamics functions. For more detailed description we refer to Nikulin et al. (2023).
@@ 3 The Missing Piece For In-Context RL
Offline RL Datasets. The datasets in this category can be considered classical, as some of them exist for more than four years now (Fu et al., 2020). They were initially proposed for offline RL, containing simple tasks with a flat structure, e.g. perform locomotion with different robots (Lu et al., 2023) or path finding in a maze. Some of them also contain data from robotic manipulators (Fu et al., 2020; Rafailov et al., 2023), or even Atari frames (Gulcehre et al., 2021). Another datasets collect data for more sophisticated environments, such as NetHack Learning Environment (Hambro et al., 2023; Kurenkov et al., 2024). However, the aforementioned datasets offer $<100$ different tasks with a fixed policy (except for the -replay datasets, which have limited coverage of various policies). This limitation makes it difficult for in-context RL to emerge from such data. To overcome this pitfall, we collected almost $30,000$ tasks with a deep ruleset structure, that are a challenging problem to solve.
Another large dataset, Open X-Embodiment (Padalkar et al., 2023), is a combination of more than 60 datasets from different robotics research labs. It consists of 527 different tasks in robotics with the demonstrations from mostly from human experts. Despite the large quantity of transitions in these datasets, they do not contain learning histories with improving policies, making their application for in-context RL quite challenging. On the contrary, our XLand-100B dataset consists of $100$ billions of transitions of RL agents’ learning histories, making it possible for in-context abilities to emerge.
@@ 4 XLand-100B Dataset
We present XLand-100B, a large dataset for in-context RL, and its smaller and simpler version XLand-Trivial-20B for faster experimentation. Together they contain about 3.5B episodes, 130B transitions and 40,000 unique tasks (see Table 2 for detailed statistics). Datasets are hosted on public S3 bucket and freely available for everyone under CC BY-SA 4.0 licence. The code is open-source and available under Apache 2.0 licence at https://github.com/dunno-lab/xland-minigrid-datasets. Next, we describe the data format, collection and evaluation.
@@ 4.1 Data Format
Storage format. We chose to store the datasets in HDF5111https://github.com/HDFGroup/hdf5 file format based on its popularity and convenience. It allows to work with large amounts of structured data without loading it to memory, store arbitrary metadata, and customise compression and chunk size to maximize the sampling throughput. We used gzip compression with default compression strength of 6, which reduced dataset size from almost 5TB+ to just $\sim$600GB for our main dataset. Using a little trick described later, we were able to reduce the size even more to just 326 GB (see Table 2). However, a naive use of compression can dramatically increase batch sampling time and slow overall training time down. We tuned HDF5 cache chunk size specifically to maximize sampling throughput for large sequence lengths. After tuning, we achieved a fourfold speedup over the naive compression, and were only two times slower compared to no compression. Given that we reduced the dataset size by a factor of 15, we see this as a good trade-off, which increases the overall affordability of the dataset. See Appendix F for throughput benchmarks.
Data and metadata format. We collect complete learning histories, i.e. for each history we store all observations, actions, rewards and dones encountered during agent training in separate HDF5 groups with unique IDs per history (see Appendix D for more details). To be compatible with DPT-like methods (Lee et al., 2023), we also store expert actions for each transition (see Section 4.2). Unlike popular formats such as RLDS [^2] https://github.com/google-research/rlds and Minari [^3] https://github.com/Farama-Foundation/Minari, we store transitions as one sequential array per history per modality. The rationale here is that under compression, it is much cheaper to sample slices of long episodes in sequence rather than sampling across different groups.
@@ 4.2 Data Collection
Figure 3: Single-task evaluation curves on 36 hard tasks for policies trained from scratch or fine-tuned from multi-task pre-trained checkpoints. See Appendix G for curves on tasks of all difficulty.
@@
At a high level, data collection was organised into three stages, namely multi-task task-conditioned pre-training; single-task fine-tuning to collect learning histories; and finally post-processing and filtering. Although we used highly optimised GPU-accelerated implementations of the base RL algorithm and environment, it still took $50,000$ GPU hours to collect the full dataset. Collecting a dataset of this size for any other environment suitable for in-context RL, e.g. Meta-World (Yu et al., 2020), would take much longer, which is unlikely to be feasible for most practitioners (Nikulin et al., 2023). Next, we describe the collection process, including the selection of the base RL algorithm and all subsequent steps. We provide exact hyperparameters for each stage in Appendix E.
@@
Pretraining. For our main XLand-100B dataset we uniformly sampled tasks from medium-1m benchmark from XLand-MiniGrid. It contains tasks of various difficulty, ranging from zero to nine rules. Unfortunately, on many hard tasks our base algorithm could not manage to converge in the time budget allocated for a single training run. On the hardest tasks, it was not even possible to get a non-zero reward at all, due to the exploration challenge that such tasks poses. In order to speed up convergence and exploration on harder tasks, we pre-train an agent in a multi-task task-conditioned manner. We expose the ruleset specification, which is usually hidden from the agent, and encode the goal and rules via embeddings, concatenating resulting encodings and passing it as an additional input to the agent. After that, we train the agent on 65k tasks simultaneously for 25B transitions. As Figure 2 shows, such an agent learns to generalize zero-shot on new tasks quite well, although we do not aim to push it to the limit, as zero-shot generalization does not produce a smooth learning history during fine-tuning. We skip this stage for XLand-Trivial-20B dataset due to the simplicity of the tasks in trivial-1m benchmark.
Finetuning. This is a key stage in the data collection process, during which we finetune a pretrained agent while recording the transitions encountered into the dataset. We finetune the agent using 8192 parallel environments for 1B transitions on 30k uniformly sampled tasks from medium-1m benchmark. We mask out the task-conditioning encoding to prevent zero-shot generalization. We record transitions only from first 32 parallel environments. This way, we can keep the size of the dataset manageable, still leaving the possibility to study scaling laws and generalization in a controlled manner. For example, we can train AD (Laskin et al., 2022) on all 30k tasks using one history per task or on $\sim$ 900 tasks using all histories per task to equalize the number of training tokens. For the XLand-Trivial-20B dataset, instead of fine-tuning, we train the agent from scratch on 10k uniformly sampled tasks from the trivial-1m benchmark, keeping all other hyperparameters the same. In the Figure 3 we show the effect of finetuning on hard tasks (with more than seven rules) compared to training from scratch. It can be seen that we are able to show strong performance even on the hardest tasks, increasing the diversity and coverage of the resulting dataset. For the same results on tasks of all levels of difficulty, see Appendix G.
@@
Postprocessing. After fine-tuning, it was necessary to additionally label the transitions with the expert actions to support DPT-like methods (Lee et al., 2023). To do this, we walk through the entire learning history with the final policy, starting from the initial hidden state for the RNN. We discuss the validity of such a labelling scheme later in the Section 4.3. Finally, all individual learning histories from different tasks were combined into one large dataset. To ensure quality, we filtered out any task with a final return below $0.3$ as an unrepresentative learning history. There were some failures, such as GPU crashes, which are inevitable during large-scale training. So any runs with corrupted data were also filtered out. In total, we filtered out about 1k tasks, leaving almost 29k tasks in the final dataset. We provide detailed statistics for each dataset in Table 2 and the final distribution of tasks by number of rules in Figure 4.
@@ 4.3 Data Evaluation
In this section we validate that the resulting XLand-100B dataset actually fulfils the two most important requirements for in-context RL, namely it contains learning histories with distinct policy improvement pattern and has expert actions for each transition (see Section 3 for a discussion). We provide analogous results for XLand-Trivial-20B in the Appendix H.
Improvement history. In the Figure 5 we show the averaged return from the learning histories separated by the number of rules. In order to better show the speed of learning on the same scale, we have normalized the x-axis for each level of difficulty, as the number of episodes can vary greatly (as it takes more time to solve complex tasks). One can see that the dataset provides a whole range of learning speeds, from very fast on easy problems to much slower on the hardest, which may be important for methods based on AD (Zisman et al., 2023; Shi et al., 2024). For a learning history averaged over full dataset see Appendix H.
@@ 5 Experiments
In this section, we investigate whether our datasets can enable an in-context RL ability. Additionally, we demonstrate how well current in-context algorithms perform across different task complexities and outline their current limitations. We take AD (Laskin et al., 2022) and DPT (Lee et al., 2023) for our experiments, the exact implementations details are in Appendix I and Appendix J. Both methods were trained on XLand-Trivial-20B and XLand-100B with $\{512,1024,2048,4096\}$ and $\{1024,2048,4096\}$ context lengths respectively. We do not include a 512 context length for the -100B dataset as we consider it too short, given the nearly 3x increase in median episode length between the two datasets. For evaluation, we run three models on $1024$ unseen tasks for $500$ episodes.
@@
DPT. Decision-Pretrained Transformer (Lee et al., 2023) is another method that exhibit in-context RL capabilities. However, in our experiments we were unable to train it so that these abilities emerge. Figure 18 demonstrates the lack of performance even on the simplest trivial tasks. We believe this is closely connected to the inability of DPT to reason in POMDP environments. For a detailed explanation, we refer the reader to Section J.1.

### v2

@@ Abstract
Following the success of the in-context learning paradigm in large-scale language and computer vision models, the recently emerging field of in-context reinforcement learning is experiencing a rapid growth. However, its development has been held back by the lack of challenging benchmarks, as all the experiments have been carried out in simple environments and on small-scale datasets. We present XLand-100B, a large-scale dataset for in-context reinforcement learning based on the XLand-MiniGrid environment, as a first step to alleviate this problem. It contains complete learning histories for nearly $30,000$ different tasks, covering $100$B transitions and $2.5$B episodes. It took $50,000$ GPU hours to collect the dataset, which is beyond the reach of most academic labs. Along with the dataset, we provide the utilities to reproduce or expand it even further. We also benchmark common in-context RL baselines and show that they struggle to generalize to novel and diverse tasks. With this substantial effort, we aim to democratize research in the rapidly growing field of in-context reinforcement learning and provide a solid foundation for further scaling.
@@ 1 Introduction
However, despite the rapid adoption of the transformer architecture in reinforcement learning (RL) after the release of Decision Transformer (DT) (Chen et al., 2021; Hu et al., 2022; Agarwal et al., 2023; Li et al., 2023), models with in-context learning capabilities appeared only recently. This delay is caused by a number of reasons. Firstly, to transition from in-weights to in-context learning, a model should be trained on tens of thousands of unique tasks (Kirsch et al., 2022). Unfortunately, even the largest RL datasets currently contain only hundreds of tasks (Padalkar et al., 2023). Secondly, it was necessary to determine the right way to provide a context to a transformer and develop the data collection pipeline, which for many methods (Laskin et al., 2022; Lee et al., 2023; Shi et al., 2024) was different from what is commonly available in existing datasets (see Section 3).
Due to the lack of suitable datasets and the high cost of collecting data in existing environments, the recent wave of in-context RL research (Laskin et al., 2022; Lee et al., 2023; Norman & Clune, 2023; Kirsch et al., 2023; Sinii et al., 2023; Zisman et al., 2023) used environments with very simple task distributions, where it was feasible to collect datasets with hundreds of tasks. While these benchmarks are affordable, they are not suitable for the comparison of methods at scale on tasks with high diversity and difficulty, which is essential for real-world applications. Because of this, the development of in-context RL is currently hindered by these factors. We believe it is crucial to address these barriers, given the essential role of in-context learning in the path to foundation models and truly generalist agents (Team et al., 2021; 2023; Kirsch et al., 2023; Lu et al., 2024; Liu et al., 2024).
@@
Data # tasks # transitions # episodes Size Open-Source Enables ICRL XLand-100B (ours) 28,876 100B 2.5B 320GB JAT 157 323M N/A 1TB GATO 596 1.5T 63M N/A Open X-Embodiment 527 N/A 2.4M 9TB AlphaStar Unplugged 1 21B 2.8M N/A ? NetHack 1 3.5B 110K 97 GB D4RL 11 N/A 40 M 4.8GB V-D4RL 4 2.4M N/A 17GB D5RL 50 N/A 2500 N/A RL Unplugged 90 80M N/A 54.1TB Procgen 16 37M N/A 500GB
@@ 2.2 XLand-MiniGrid
Figure 1: Visualization of a generic XLand-MiniGrid environment. Grid layout should be selected in advance, while the positions of the objects are randomized on each reset. For the dataset we use simpler layout with just one room, see Appendix M.
@@
Environment. XLand-MiniGrid is a complete rewrite of MiniGrid (Chevalier-Boisvert et al., 2023) in JAX (Bradbury et al., 2018), incorporating a notion of rules and goals from XLand (Team et al., 2023). Leveraging JAX, it can run on a GPU or TPU accelerators at millions of steps per seconds. At its core, it is a goal-oriented grid-world environment with simple underlying dynamics, partial observability and sparse rewards. The action space is simple, consisting mainly of navigation and interaction with game objects, such as opening a door or picking and placing items. Observations are symbolic "images" encoding the agent surrounding as tile and color ID’s. Rules are functions that can change the state of the environment based on some conditions, e.g. when two specific objects are places near each other, they both disappear and one new object is placed. Goals are similar, except they only validate some predefined conditions and do not change anything. Composing different rules and goals together we can create new tasks with varying reward and dynamics functions. For more detailed description we refer to Nikulin et al. (2023).
@@ 3 The Missing Piece For In-Context RL
Offline RL Datasets. The datasets in this category can be considered classical, as some of them exist for more than four years now (Fu et al., 2020). They were initially proposed for offline RL, containing simple tasks with a flat structure, e.g. perform locomotion with different robots (Lu et al., 2023) or path finding in a maze. Some of them also contain data from robotic manipulators (Fu et al., 2020; Rafailov et al., 2023), or even Atari frames (Gulcehre et al., 2021). Other datasets collect data for more sophisticated environments, such as NetHack Learning Environment (Hambro et al., 2023; Kurenkov et al., 2024) or ProcGen (Cobbe et al., 2019; Mediratta et al., 2024). However, the aforementioned datasets offer $<100$ different tasks with a fixed policy (except for the -replay datasets, which have limited coverage of various policies). This limitation makes it difficult for in-context RL to emerge from such data. To overcome this pitfall, we collected almost $30,000$ tasks with a deep ruleset structure, that are a challenging problem to solve.
Another large dataset, Open X-Embodiment (Padalkar et al., 2023), is a combination of more than 60 datasets from different robotics research labs. It consists of 527 different tasks in robotics with the demonstrations from mostly human experts. Despite the large quantity of transitions in these datasets, they do not contain learning histories with improving policies, making their application for in-context RL quite challenging. On the contrary, our XLand-100B dataset consists of $100$ billions of transitions of RL agents’ learning histories, making it possible for in-context abilities to emerge.
@@ 4 XLand-100B Dataset
We present XLand-100B, a large dataset for in-context RL, and its smaller and simpler version XLand-Trivial-20B for faster experimentation. Together they contain about 3.5B episodes, 130B transitions and 40,000 unique tasks (see Table 2 for detailed statistics). Datasets are hosted on public S3 bucket and will be freely available for everyone under CC BY-SA 4.0 licence. Next, we describe the data format, collection and evaluation.
@@ 4.1 Data Format
Storage format. We chose to store the datasets in HDF5111https://github.com/HDFGroup/hdf5 file format based on its popularity and convenience. It allows to work with large amounts of structured data without loading it to memory, store arbitrary metadata, and customise compression and chunk size to maximize the sampling throughput. We used gzip compression with default compression strength of 6, which reduced dataset size from almost 5TB+ to just $\sim$600GB for our main dataset. Using a little trick described later, we were able to reduce the size even more to just 326 GB (see Table 2). However, a naive use of compression can dramatically increase batch sampling time and slow overall training time down. We tuned HDF5 cache chunk size specifically to maximize sampling throughput for large sequence lengths. After tuning, we achieved a fourfold speedup over the naive compression, and were only two times slower compared to no compression. Given that we reduced the dataset size by a factor of 15, we see this as a good trade-off, which increases the overall affordability of the dataset. See Appendix C for throughput benchmarks.
Data and metadata format. We collect complete learning histories, i.e. for each history we store all observations, actions, rewards and dones encountered during agent training in separate HDF5 groups with unique IDs per history (see Appendix B for more details). To be compatible with DPT-like methods (Lee et al., 2023), we also store expert actions for each transition (see Section 4.2). Unlike popular formats such as RLDS [^2] https://github.com/google-research/rlds and Minari [^3] https://github.com/Farama-Foundation/Minari, we store transitions as one sequential array per history per modality. The rationale here is that under compression, it is much cheaper to sample slices of long episodes in sequence rather than sampling across different groups.
@@ 4.2 Data Collection
Figure 3: Single-task evaluation curves on 36 hard tasks for policies trained from scratch or fine-tuned from multi-task pre-trained checkpoints. See Appendix D for curves on tasks of all difficulty.
@@
At a high level, data collection was organized into three stages, namely multi-task task-conditioned pre-training; single-task fine-tuning to collect learning histories; and finally post-processing and filtering. Although we used highly optimised GPU-accelerated implementations of the base RL algorithm and environment, it still took $50,000$ GPU hours to collect the full dataset. Collecting a dataset of this size for any other environment suitable for in-context RL, e.g. Meta-World (Yu et al., 2020), would take much longer, which is unlikely to be feasible for most practitioners (Nikulin et al., 2023). Next, we describe the collection process, including the selection of the base RL algorithm and all subsequent steps. We provide exact hyperparameters for each stage in Appendix O.
@@
Pretraining. For our main XLand-100B dataset we uniformly sampled tasks from medium-1m benchmark from XLand-MiniGrid. It contains tasks of various difficulty, ranging from zero to nine rules. Unfortunately, on many hard tasks our base algorithm could not manage to converge in the time budget allocated for a single training run. On the hardest tasks, it was not even possible to get a non-zero reward at all, due to the exploration challenge that such tasks poses. In order to speed up convergence and exploration on harder tasks, we pre-train an agent in a multi-task task-conditioned manner and also use simpler grid layout with just one room (see Appendix M). We expose the ruleset specification, which is usually hidden from the agent, and encode the goal and rules via embeddings, concatenating resulting encodings and passing it as an additional input to the agent. After that, we train the agent on 65k tasks simultaneously for 25B transitions. As Figure 2 shows, such an agent learns to generalize zero-shot on new tasks quite well, although we do not aim to push it to the limit, as zero-shot generalization does not produce a smooth learning history during fine-tuning. We skip this stage for XLand-Trivial-20B dataset due to the simplicity of the tasks in trivial-1m benchmark.
Finetuning. This is a key stage in the data collection process, during which we finetune a pretrained agent while recording the transitions encountered into the dataset. We finetune the agent using 8192 parallel environments for 1B transitions on 30k uniformly sampled tasks from medium-1m benchmark. We mask out the task-conditioning encoding to prevent zero-shot generalization. We record transitions only from first 32 parallel environments. This way, we can keep the size of the dataset manageable, still leaving the possibility to study scaling laws and generalization in a controlled manner. For example, we can train AD (Laskin et al., 2022) on all 30k tasks using one history per task or on $\sim$ 900 tasks using all histories per task to equalize the number of training tokens. For the XLand-Trivial-20B dataset, instead of fine-tuning, we train the agent from scratch on 10k uniformly sampled tasks from the trivial-1m benchmark, keeping all other hyperparameters the same. In the Figure 3 we show the effect of finetuning on hard tasks (with more than seven rules) compared to training from scratch. It can be seen that we are able to show strong performance even on the hardest tasks, increasing the diversity and coverage of the resulting dataset. For the same results on tasks of all levels of difficulty, see Appendix D.
@@
Postprocessing. After fine-tuning, it was necessary to additionally label the transitions with the expert actions to support DPT-like methods (Lee et al., 2023). To do this, we walk through the entire learning history with the final policy, starting from the initial hidden state for the RNN. We evaluate the validity of such a labelling scheme later in the Section 4.3. Finally, all individual learning histories from different tasks were combined into one large dataset. To ensure quality, we filtered out any task with a final return below $0.3$ as an unrepresentative learning history. There were some failures, such as GPU crashes, which are inevitable during large-scale training. So any runs with corrupted data were also filtered out. In total, we filtered out about 1k tasks, leaving almost 29k tasks in the final dataset. We provide detailed statistics for each dataset in Table 2 and the final distribution of tasks by number of rules in Figure 4.
@@ 4.3 Data Evaluation
In this section we validate that the resulting XLand-100B dataset actually fulfils the two most important requirements for in-context RL, namely it contains learning histories with distinct policy improvement pattern and has expert actions for each transition (see Section 3 for a discussion). We provide analogous results for XLand-Trivial-20B in the Appendix E.
Figure 6: Agreement between actions predicted by the expert and the actual actions in the learning history. We use final PPO policy as an expert for actions labeling. As can be seen agreement is growing towards the end of learning history.
Improvement history. In the Figure 5 we show the averaged return from the learning histories separated by the number of rules. In order to better show the speed of learning on the same scale, we have normalized the x-axis for each level of difficulty, as the number of episodes can vary greatly (as it takes more time to solve complex tasks). One can see that the dataset provides a whole range of learning speeds, from very fast on easy problems to much slower on the hardest, which may be important for methods based on AD (Zisman et al., 2023; Shi et al., 2024). For a learning history averaged over full dataset see Appendix E.
@@
5 Experiments
In this section, we investigate whether our datasets can enable an in-context RL ability. Additionally, we demonstrate how well current in-context algorithms perform across different task complexities and outline their current limitations. We take AD (Laskin et al., 2022) and DPT (Lee et al., 2023) for our experiments, the exact implementations details are in Appendix F and Appendix G. Both methods were trained on XLand-Trivial-20B and XLand-100B with $\{512,1024,2048,4096\}$ and $\{1024,2048,4096\}$ context lengths respectively. We do not include a 512 context length for the -100B dataset as we consider it too short, given the nearly 3x increase in median episode length between the two datasets. For evaluation, we run three models on $1024$ unseen tasks for $500$ episodes.
@@ 5 Experiments
DPT. Decision-Pretrained Transformer (Lee et al., 2023) is another method that exhibit in-context RL capabilities. However, in our experiments we were unable to train it so that these abilities emerge. Figure 18 in Appendix K demonstrates the lack of performance even on the simplest trivial tasks. We believe this is closely connected to the inability of DPT to reason in POMDP environments. For a detailed investigation, we refer the reader to Appendix H.